OpenStudy (anonymous):

The sum of 9/26 and -4/13 is =?

OpenStudy (nurali):

\[\frac{ 9 }{ 26 }-\frac{ 4 }{ 13 }\] \[\frac{ 13*9-4*26 }{ 26*13 }\] \[\frac{ 1 }{ 26 }\]